Main INDEX
Monthly INDEX
PREV
NEXT
User name vasulk
Log entry time 13:36:37 on January19,2003
Entry number 92536
This entry is a followup to: 92534
keyword=Transverse lasers (better format)
When I came to the Counting House at 11:15 am to reboot the spectrum analyzer computer, I noticed that the pumping chamber RTD's were reading as if the lasers were not on. At closer inspection, the transverse laser GUI's showed that all the lasers were set to 15 amps. Since this morning was the first time we went to 270 degrees at 5 GeV, the problem had not been noticed before. Looking at the end-of-run Halog entries, I found that the readback for the lasers was fine up until 12:30 pm Friday, January 17, 2003. But the next end of run wasn't until 19:30 that night, and at that time the lasers were reading back 15 amps.
We quickly rotated the field to 0 degrees to polarize the cell, while I reset the currents for the transverse lasers. Once at 0 degrees, we performed an NMR and EPR. The NMR was 32.5% and EPR gave 32.0 %. The EPR spectrum looked good, and no masing was seen.
When target operators change the target holding field and change the laser configuration, they should keep an eye on the pumping chamber RTD's (numbers 6 and 7) for at least 15 minutes to make sure things really are fine. With the longitudinal lasers on, RTD # 7 should be stable at 195 degrees C, and RTD # 6 should be ~ 175 degrees. With the transverse lasers on, RTD # 7 should be stable at 187 degrees C, and RTD # 6 should be ~ 169 degrees. If there is a problem page one of the target experts.